OpenStudy (anonymous):

If the circumference is numerically equal to 3 times the area of circle , then find the radius of the circle.

OpenStudy (dumbcow):

oops i didn;t read the question right

OpenStudy (anonymous):

can any one explain it also ????

OpenStudy (dumbcow):

2pi*r = 3pi*r^2 divide by pi*r 2 = 3r r = 2/3

OpenStudy (anonymous):

thnx dude !!!!!

OpenStudy (anonymous):

okay. 2pi*r=3*pi*r^2 2pi=3pi*r r=2pi/3pi r=2/3=.6667

OpenStudy (anonymous):

heyyy but the answer is given 6 !!!!